OpenStudy (anonymous):

WILL GIVE MEDAL<<<<<<<<<<<<<<< Angle H and angle J are complementary. Angle H measures x+5 and angle J measures 2x-5 find the measure of angle H

OpenStudy (anonymous):

easy x+5+2x-5=90 3x=90 x=30 x+5=35 2x-5=55